A charged Q of 2.5uC is in two dimensional electric field with horizontal and vertical components of 35.0 N/C and 52.0 N/C, respectively. Find the magnitude and direction of the electric force on charge Q.


1
Expert's answer
2022-03-18T12:01:41-0400

Now we know that

"E_{net}=\sqrt{E_1+E_2}"


"E_{net}=\sqrt{35^2+52^2}=\sqrt{1225+2704}=\sqrt{3929}=62.68N/C"

Force

"F=qE"


"F=2.5\times10^{-6}\times62.68=1.56\times10^{-4}N"

Direction


"\phi =tan^{-1}(\frac{52}{35})=tan^{-1}(1.48)=56°"